# Geometric Scaling of Bayesian Inference in LLMs

### v1

@@ 1. Introduction
Large language models have achieved striking performance across natural language, coding, mathematics, and reasoning tasks (brown2020language; chen2021evaluating; openai2024reasoning). Yet their internal computations remain only partially understood. A central question is whether transformers merely approximate statistical associations at scale, or whether they implement more principled forms of probabilistic inference.
@@
Recent work (Paper 1 of this trilogy (concurrent_bayes)) demonstrated that small transformers trained in controlled “Bayesian wind tunnels” perform exact Bayesian inference on bijection-learning and HMM-filtering tasks: posterior entropy, KL divergence, and predictive distributions match analytic solutions to within 0.1 bits. Paper 2 showed that these behaviors arise from specific geometric mechanisms created by gradient dynamics: value manifolds ordered by entropy, orthogonal key frames defining hypothesis directions, and layerwise attention sharpening implementing a geometric Bayes rule.
@@
Papers 1 and 2 in this series established that transformers can implement exact Bayesian inference in controlled “wind—-tunnel” settings, and that gradient dynamics generically sculpt the value and key spaces into geometric substrates that support such inference. However, these results were obtained in synthetic domains with analytically specified likelihoods. The present work answers a distinct question: do production-grade LLMs, trained on heterogeneous natural language, spontaneously develop the same geometric signatures, and do these signatures track evidence integration during inference? We show that the three hallmarks of Bayesian geometry—-low-dimensional value manifolds, orthogonal hypothesis frames, and evidence-dependent movement along entropy-aligned directions—-persist across four model families (Pythia, Phi-2, Llama-3, Mistral). Moreover, we provide the first large-scale evidence that these structures are functionally engaged during inference in a naturalistic task (SULA), even when their causal role is distributed rather than bottlenecked. This establishes that Bayesian geometry is not an artifact of synthetic tasks, but a stable inductive bias of modern transformers.
@@ 4.1. Model Selection
Pythia-410M (biderman2023pythia): GPT-NeoX architecture, 24 layers, 16 attention heads, 1024 hidden dimensions. Trained on the Pile corpus (800GB diverse text) with full training transparency. Represents canonical standard MHA trained on general-purpose data.
@@ Conclusion.
The Mistral family should therefore be viewed not as a counterexample to Bayesian geometry but as a boundary case that reveals how architectural constraints selectively modulate dynamic Bayesian computation. Static geometric structure—entropy-ordered manifolds and hypothesis frames—persists across all Mistral variants, while dynamic refinement is diminished by local attention and MoE routing. This provides a natural explanation for the observed behavior and connects directly to the theoretical predictions of Paper 2.
@@ 7. Discussion
Our analysis complements circuit-level studies such as induction heads, copy heads, and pattern-matching mechanisms (elhage2021mathematical; olsson2022context). Those works identify specialized components; our findings describe the global geometric scaffold in which such components operate. Attention focusing determines which tokens are consulted; key orthogonality creates separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional axes. Mapping specific heads onto regions or branches of this geometry is a natural next step.
@@ 7.2. Related Work
Tuned-lens methods (belrose2023tunedlens) decode intermediate-layer predictions by training a small linear adapter that maps hidden states back to the model’s output space. These approaches probe what the model would predict at each layer, whereas our value-manifold analysis characterizes how the model represents predictive uncertainty. The two perspectives are complementary: PC1 coordinates correlate with entropy and tuned-lens predictions, suggesting that uncertainty is encoded geometrically along a small number of directions. Establishing a principled correspondence between tuned-lens logits and manifold coordinates is an important next step.
@@
Recent work in computational mechanics (marks2024computationalmechanics) shows that belief states in small transformers can be linearly decoded from the residual stream, revealing simple geometric representations of latent uncertainty. Our results are consistent with this interpretation: production models appear to maintain analogous belief-state structure, but encoded predominantly in the value space of the final attention layer. This distinction clarifies where uncertainty lives in deeper networks and suggests that value manifolds may serve as the canonical substrate for model beliefs. Preliminary analyses indicate that coordinates along our PC1 axis correlate with tuned-lens predictions and residual-stream belief decoders, but a systematic alignment between value-manifold axes and decoded belief variables requires additional work and we leave a full treatment to future work.
@@
Studies of attention-entropy trajectories (voita2019analyzing; clark2019does) report that sharpening can be unstable or highly input-dependent. Our layerwise entropy results align with these findings: MHA models exhibit strong, stable focusing; GQA models show weaker but monotone reduction; and sliding-window and MoE architectures often display non-monotone or noisy behavior. These dynamics reflect architectural constraints on global routing rather than absence of Bayesian computation, consistent with the separation of static and dynamic geometry we observe.
@@
The constrained-belief-update model of early-layer attention (dalal2025gradient) predicts that attention patterns should stabilize early, while finer-grained posterior refinements should occur in value representations. This matches precisely the frame—-precision dissociation predicted in Paper 2 and observed empirically here: keys define a stable hypothesis frame, while values encode uncertainty refinement along low-dimensional axes.
@@
Normalization layers and architectural components influence geometric clarity. Recent analyses (ba2016layer) show that layer normalization, RoPE embeddings, and GQA induce characteristic patterns of anisotropy and dimensionality in the residual stream. Our results refine this picture by showing that such architectural choices modulate dynamic signatures (especially attention entropy reduction) while leaving the static Bayesian geometry - value manifolds and orthogonal key frames - largely intact.
@@
Circuit-level analyses (elhage2021framework; olsson2022icl) identify specific mechanisms such as induction heads, pattern-matchers, and copy circuits. Our work operates at a complementary scale: we provide a geometric account of the global representational substrate in which such circuits operate. Attention mechanisms determine which evidence is consulted; key frames define separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional coordinates. Mapping specific circuits onto these global geometric structures is a promising direction for future work.

### v2

@@ 1. Introduction
Large language models have achieved striking performance across natural language, coding, mathematics, and reasoning tasks (Brown et al., 2020; Chen et al., 2021; OpenAI, 2024). Yet their internal computations remain only partially understood. A central question is whether transformers merely approximate statistical associations at scale, or whether they implement more principled forms of probabilistic inference.
@@
Recent work (Paper 1 of this trilogy (Aggarwal et al., 2025a)) demonstrated that small transformers trained in controlled “Bayesian wind tunnels” perform exact Bayesian inference on bijection-learning and HMM-filtering tasks: posterior entropy, KL divergence, and predictive distributions match analytic solutions to within 0.1 bits. Paper 2 showed that these behaviors arise from specific geometric mechanisms created by gradient dynamics: value manifolds ordered by entropy, orthogonal key frames defining hypothesis directions, and layerwise attention sharpening implementing a geometric Bayes rule.
@@
Clarification on “Bayesian inference.”
Throughout this trilogy, “Bayesian inference” refers to the Bayesian posterior predictive over latent task variables—e.g., filtering posteriors over hidden states—not a posterior over network weights. This is a statement about the function the transformer computes, not about weight-space uncertainty.
@@
Papers 1 and 2 in this series established that transformers can implement exact Bayesian inference in controlled “wind—-tunnel” settings, and that gradient dynamics generically sculpt the value and key spaces into geometric substrates that support such inference. However, these results were obtained in synthetic domains with analytically specified likelihoods. The present work answers a distinct question: do production-grade LLMs, trained on heterogeneous natural language, spontaneously develop the same geometric signatures, and do these signatures track evidence integration during inference? We show that the three hallmarks of Bayesian geometry—-low-dimensional value manifolds, orthogonal hypothesis frames, and evidence-dependent movement along entropy-aligned directions—-persist across four model families (Pythia, Phi-2, Llama-3, Mistral). Moreover, we provide the first large-scale evidence that these structures are functionally engaged during inference in a naturalistic task (SULA), even when their causal role is distributed rather than bottlenecked. This establishes that Bayesian geometry is a stable inductive bias of modern transformers, not an artifact of synthetic tasks.
@@ 4.1. Model Selection
Pythia-410M (Biderman et al., 2023): GPT-NeoX architecture, 24 layers, 16 attention heads, 1024 hidden dimensions. Trained on the Pile corpus (800GB diverse text) with full training transparency. Represents canonical standard MHA trained on general-purpose data.
@@ Conclusion.
The Mistral family should therefore be viewed as a boundary case that reveals how architectural constraints selectively modulate dynamic Bayesian computation, rather than as a counterexample to Bayesian geometry. Static geometric structure—entropy-ordered manifolds and hypothesis frames—persists across all Mistral variants, while dynamic refinement is diminished by local attention and MoE routing. This provides a natural explanation for the observed behavior and connects directly to the theoretical predictions of Paper 2.
@@ 7. Discussion
Our analysis complements circuit-level studies such as induction heads, copy heads, and pattern-matching mechanisms (Elhage et al., 2021a; Olsson et al., 2022a). Those works identify specialized components; our findings describe the global geometric scaffold in which such components operate. Attention focusing determines which tokens are consulted; key orthogonality creates separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional axes. Mapping specific heads onto regions or branches of this geometry is a natural next step.
@@ 7.2. Related Work
Tuned-lens methods (Belrose et al., 2023) decode intermediate-layer predictions by training a small linear adapter that maps hidden states back to the model’s output space. These approaches probe what the model would predict at each layer, whereas our value-manifold analysis characterizes how the model represents predictive uncertainty. The two perspectives are complementary: PC1 coordinates correlate with entropy and tuned-lens predictions, suggesting that uncertainty is encoded geometrically along a small number of directions. Establishing a principled correspondence between tuned-lens logits and manifold coordinates is an important next step.
@@
Recent work in computational mechanics (Marks and Tegmark, 2024) shows that belief states in small transformers can be linearly decoded from the residual stream, revealing simple geometric representations of latent uncertainty. Our results are consistent with this interpretation: production models appear to maintain analogous belief-state structure, but encoded predominantly in the value space of the final attention layer. This distinction clarifies where uncertainty lives in deeper networks and suggests that value manifolds may serve as the canonical substrate for model beliefs. Preliminary analyses indicate that coordinates along our PC1 axis correlate with tuned-lens predictions and residual-stream belief decoders, but a systematic alignment between value-manifold axes and decoded belief variables requires additional work and we leave a full treatment to future work.
@@
Studies of attention-entropy trajectories (Voita et al., 2019; Clark et al., 2019) report that sharpening can be unstable or highly input-dependent. Our layerwise entropy results align with these findings: MHA models exhibit strong, stable focusing; GQA models show weaker but monotone reduction; and sliding-window and MoE architectures often display non-monotone or noisy behavior. These dynamics reflect architectural constraints on global routing rather than absence of Bayesian computation, consistent with the separation of static and dynamic geometry we observe.
@@
The constrained-belief-update model of early-layer attention (Aggarwal et al., 2025b) predicts that attention patterns should stabilize early, while finer-grained posterior refinements should occur in value representations. This matches precisely the frame—-precision dissociation predicted in Paper 2 and observed empirically here: keys define a stable hypothesis frame, while values encode uncertainty refinement along low-dimensional axes.
@@
Normalization layers and architectural components influence geometric clarity. Recent analyses (Ba et al., 2016) show that layer normalization, RoPE embeddings, and GQA induce characteristic patterns of anisotropy and dimensionality in the residual stream. Our results refine this picture by showing that such architectural choices modulate dynamic signatures (especially attention entropy reduction) while leaving the static Bayesian geometry - value manifolds and orthogonal key frames - largely intact.
@@
Circuit-level analyses (Elhage et al., 2021b; Olsson et al., 2022b) identify specific mechanisms such as induction heads, pattern-matchers, and copy circuits. Our work operates at a complementary scale: we provide a geometric account of the global representational substrate in which such circuits operate. Attention mechanisms determine which evidence is consulted; key frames define separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional coordinates. Mapping specific circuits onto these global geometric structures is a promising direction for future work.